Maxwell owes Mr Langa 5000 due in three months and 2000 due in six months. Maxwell offers to pay 3000 immediately if he can be allowed to pay the balance in one year. Mr Langa agrees on condition that an interest rate of 16% is charged on the 3000 paid now and all the loan amounts. How much will Maxwell pay at the end of the year?
Answer by Theo(13342)

if this is simple interest, then i think it can be solved in the following manner.
simple interest formula is i = p * r * n
i is the interest.
p is the principal.
r is the interest rate per time period.
n is the number of time periods.
with 5000 due in 3 months, the formula becomes i = 5000 * .16/12 * 3 = 200, and 2000 due in 6 months, becomes i = 2000 * .16/12 * 6 = 160 for total interest charge of 360.
for 3000 due now, there is no interest charge.
for the balance due in 12 months, the formula becomes i = b * .16/12 * 12.
since i must be equal to 360, then the formula becomes 360 = b * .16/12 * 12.
solve for b to get b = 360 / (.16/12 * 12) = 2250.
mister langa will get the same total interest whether maxwell pays him 5000 in 3 months and 2000 in 6 months or 3000 now and 2250 in 12 months.
i'm not 100% certain that this is the way to analyze it, but it makes sense in terms of the simple interest formula and how it' used.
here's a reference on the simple interest formula.
https://www.cuemath.com/commercial-math/simple-interest/
not that they use percent interest and then divide it by 100.
i use interest, which is percent interest already divided by 100.
16% / 100 = .16
.16 * 100 = 16%